Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Replacement in probability : concept

Expert replies
by bblast » Sun Jul 24, 2011 12:07 am
1>What is the probability that one of the two integers randomly selected from range 20-29 is prime and the other is a multiple of 3?

[spoiler]{2/10 * 3/10} *2 = .12[/spoiler]


2>A basket contains 13 fruits: 5 apples and 8 pears. What is the probability that 2 fruits drawn from the basket at random will be pears?

[spoiler]8/13*7/12 = 14/39[/spoiler]




Both question above are from gmat club tests. My query is when do we consider that we are talking about replacements ? I mean the answer trap to the second question could have been 8/13*8/13 (if we considered that the fruit is replaced after withdrawal)


In my first question if we considered that the prime is not being put back then the solution would be {2/10 * 3/9} *2
Cheers !!

Quant 47-Striving for 50
Verbal 34-Striving for 40

My gmat journey :
https://www.beatthegmat.com/710-bblast-s ... 90735.html
My take on the GMAT RC :
https://www.beatthegmat.com/ways-to-bbla ... 90808.html
How to prepare before your MBA:
https://www.youtube.com/watch?v=upz46D7 ... TWBZF14TKW_
Join the discussion
Source: — Problem Solving |

by knight247 » Sun Jul 24, 2011 1:02 am
We consider replacement only and only when it is stated or implicitly understood. In the first question we consider it as replacement simply because you picking a number from there does not deplete one number from the pool. But on the other hand if you had the numbers 20-29 written on chits of paper and placed in a bowl and 2 chits were picked out without replacement that would deplete numbers from the bowl. And in the second scenario the answer would be {2/10 * 3/9} *2. This concept is also applied to question 2


In question 2. Thinking outside the scope of mathematics....When a fruit is taken out of a basket the same fruit cannot be picked out of the basket again. You can only select from one of the remaining fruits. So taking a fruit out of the basket reduces the number of possible options by one.
Last edited by knight247 on Sun Jul 24, 2011 3:23 am, edited 1 time in total.
Join the discussion

by naveen451 » Sun Jul 24, 2011 2:37 am
in first question
{2/10*3/10}*2 is the ans

whatis the need of *2
Join the discussion

by knight247 » Sun Jul 24, 2011 3:08 am
Specific order is not mentioned in the problem. It could be P(Prime)*P(# divisible by3)+P(# divisible by3)*P(Prime). Hence the 2
Join the discussion